OpenStudy (anonymous):

What is the inverse function of 2 + the square root of 5x-3?

OpenStudy (anonymous):

y = 2 + √(5x-3) First step of finding inverse: swap x's and y's x = 2 + √(5y-3) Put in terms of y x - 2 = √(5y-3) (x-2)^2 = 5y-3 x^2-4x+4 = 5y-3 x^2-4x+7=5y y = (x^2 - 4x + 7)/5 Is it a function? Yes So y = (x^2 - 4x + 7)/5 is the inverse function

OpenStudy (anonymous):

Thank you, the constant in front of the square root confused me.